given the measures of three angles of a triangle .explain how to find the ratio of the lengths of two sides of the triangle.
(measures are not given ,we have to explain the method.question is from law of cosines.)










Found 2 solutions by rothauserc, Alan3354:
Answer by rothauserc(4718)

I would use the law of sines
:
a/sin(A) = b/sin(B)
:
a and b are the lengths of two sides and A, B are the measures in degrees of the two angles opposite a, b
:
a/sin(A) = b/sin(B)
:
cross multiply the two fractions
:
a * sin(B) = b * sin(A)
:
sin(A)/sin(B) = a/b
:

Answer by Alan3354(69443)

given the measures of three angles of a triangle .explain how to find the ratio of the lengths of two sides of the triangle.
(measures are not given ,we have to explain the method.question is from law of cosines.)
====================
The Cosine Law can be used for this, but it's tedious.
---
Use the Law of Sines:
a/sin(A) = b/sin(B) = c/sin(C)